OpenStudy (anonymous):

force in field is phi(x,y,z) has direction gradient of phi . calculate grad phi phi(x,y,z)=((x-1)^2) - ((y+1)^2) force in field is phi(x,y,z) has direction gradient of phi . calculate grad phi phi(x,y,z)=((x-1)^2) - ((y+1)^2) @Mathematics

OpenStudy (amistre64):

grad phi is just the partial differentials right?

OpenStudy (amistre64):

x' = 2(x+1) y' = 2(2+1) z' = 0